# Dissociation driven quarkonium spin alignment in Pb–Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV

### v1

@@ II.2.4 Gluonic Dissociation
The gluon-induced dissociation (or gluonic dissociation) of quarkonia in the QGP medium is another dissociation mechanism. This process arises when a thermal gluon interacts inelastically with a color-singlet quarkonium state, making a transition into a color-octet state, which subsequently breaks apart. The corresponding dissociation cross-section depends on the gluon energy spectrum and temperature profile of the medium. Following the formalism introduced by Peskin and Bhanot [47, 48] and extended to finite-temperature QCD [49, 50, 30, 31], the gluon-dissociation cross section for a quarkonium state is given by,
$\sigma_{\rm d,nl}(E_{g})=\frac{\pi^{2}\alpha_{s}^{u}E_{g}}{N_{c}^{2}}\sqrt{\frac{{m_{q}}}{E_{g}+E_{nl}}}\\ \;\;\;\;\;\times\left(\frac{l|J_{nl}^{q,l-1}|^{2}+(l+1)|J_{nl}^{q,l+1}|^{2}}{2l+1}\right)$ (23)
where, $m_{q}$ is the mass of the charm/bottom quark, $E_{g}$ is the incident gluon energy, and $E_{nl}$ denotes the energy eigenvalue of the quarkonium state. The coupling constant $\alpha_{s}^{u}$ is scaled as $\alpha_{s}^{u}=\alpha_{s}(\alpha_{s}m_{c}^{2}/2)$. The $g_{nl}(r)$ and $h_{ql^{\prime}}(r)$ are the color-singlet and color-octet wave functions with orbital angular momenta $l$ and $l^{\prime}$, respectively. The octet wave function $h_{\rm ql^{\prime}}(r)$ has been obtained by solving the Schrödinger equation with the octet potential $V_{8}=\alpha_{\rm eff}/8r$. The value of $q$ is determined by using the conservation of energy, $q=\sqrt{m_{c}(E_{g}+E_{nl})}$.
The probability density $J_{nl}^{ql^{{}^{\prime}}}$ encodes the overlap between the initial bound and final continuum states and is computed using the in-medium singlet and octet wave functions,
$J_{\rm nl}^{\rm ql^{\prime}}=\int_{0}^{\infty}dr\;r\;g^{*}_{nl}(r)\;h_{ql^{\prime}}(r)$ (24)
@@
This modified distribution plays an essential role in determining the gluonic dissociation rate and, consequently, the momentum-dependent dissociation of quarkonia in the QGP. The thermal averaged gluonic dissociation, $\Gamma_{\rm gd,nl}$, can then be obtained by folding the cross section with this modified Bose–Einstein distribution of thermal gluons in the particle rest frame [30, 31], as given by;
$\Gamma_{\rm gd,nl}(\tau,p_{\rm T},b)=\frac{g_{d}}{4\pi^{2}}\int_{0}^{\infty}\int_{0}^{\pi}\frac{dp_{g}\,d\theta\,\sin\theta\,p_{g}^{2}\sigma_{d,nl}(E_{g})}{e^{\{\frac{\gamma E_{g}}{T_{\rm eff}}(1+v_{Q}\cos\theta)\}}-1}$ (26)
@@
Furthermore, we have obtained the net decay width $\Gamma_{\rm D}$ for quarkonia in-medium via taking the sum over Eq. 22, and  26, given as;
$\Gamma_{\rm D,nl}=\Gamma_{\rm damp,nl}+\Gamma_{\rm gd,nl}$ (27)
@@ II.3 Spin alignment of quarkonia
The dissociation probability of $i^{\rm th}$ spin state in the presence of rotation is given by;
$f_{i}=\exp\left[{-\mathop{\text{\large$\int_{\text{\normalsize$\scriptstyle\kern-1.22911pt\tau_{0}$}}^{\text{\normalsize$\scriptstyle\tau_{f}$}}$}}\nolimits\Gamma_{\rm D,nl}(\tau,p_{\rm T})\;d\tau}\right].$ (28)
Here $i$ takes on the values 0,+1,10,+1,-1, representing the three distinct spin projections corresponding to the states 0, ++1, and 1-1, respectively. The $\tau_{0}$ and $\tau_{f}$ are the proper times corresponding to the initial temperature $T_{0}$ and freeze-out temperature $T_{f}$ for the QGP medium. The dissociation rate, denoted as $\Gamma_{\rm D,nl}$, is intricately linked to both the rotational dynamics and the specific spin projection states, $i$.
The spin alignment of vector mesons is measured via the elements of the spin density matrix $\rho$. For vector meson, the $00^{\rm th}$ element of the $3\times 3$ spin density matrix is defined in terms of the $\Gamma_{\rm D,nl}$;
$\rho_{00}=\frac{f_{0}}{f_{+}+f_{0}+f_{-}}$ (29)
@@ III Results and Discussions
Quarkonium suppression and spin alignment provide complementary insights into the interaction between heavy quark–antiquark pairs and the quark–gluon plasma. Following that, we evaluated how different dissociation processes modify the spin alignment factor $\rho_{00}$ for the $J/\psi$, $\psi$(2S), $\Upsilon$(1S), and $\Upsilon$(2S) states, as defined in Eq. 29. Here, results are obtained by numerically solving the Schrödinger equation, which provides temperature, vorticity, and spin-projection-dependent eigenvalues and wave functions. The imaginary component of the potential, together with the quarkonium wave function, determines the collisional damping width, whereas the binding energy inferred from the eigenvalue governs the gluon-dissociation width. As reported in the earlier findings [33, 32], the collisional damping mechanism dominates over gluonic dissociation across the thermal evolution of the medium.
@@
The decay width illustrated in Fig. 2 shows the $p_{\rm T}$ dependence for $J/\psi$ (upper left), $\psi$(2S) (upper right), $\Upsilon$(1S) (lower left), and $\Upsilon$(2S) (lower right) in Pb–Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV in the absence in-medium rotation ($C=0$ fm) at $T$ = 175, 300, and 400 MeV. The results indicate a strong sensitivity of the dissociation rate to both temperature and momentum. At $T$ = 400 MeV, the $J/\psi$ exhibits an enhanced dissociation rate, which gradually decreases with increasing $p_{\rm T}$. The sudden drop in $\Gamma_{\rm D}$ around $p_{\rm T}\sim$ 10 GeV/c reflects that $T_{\rm eff}<<T=400$ MeV. Further, at $T$ = 300, and 175 MeV, the effect of $T_{\rm eff}$ significantly diminishes and $\Gamma_{\rm D}$ decreases as a function of $p_{\rm T}$. A similar trend is evident for other quarkonium states, including $\psi$(2S), $\Upsilon$(1S), and $\Upsilon$(2S). In particular, at temperatures of 300 and 175 MeV, $\psi$(2S) exhibits a higher dissociation probability than $J/\psi$ across the entire $p_{\rm T}$ range. This can be attributed to the comparatively lower binding energy of $\psi$(2S) relative to $J/\psi$. Consequently, the reduced survival probability of $\psi$(2S) further supports the sequential melting scenario in the QGP medium. Furthermore, due to the higher mass of $\Upsilon$(1S) and $\Upsilon$(2S) states experience relatively stable effective temperatures as they traverse through the medium, as illustrated in Fig. 1. As a result, the dissociation probability for $\Upsilon$(1S) exhibits a modest decreasing trend towards higher $p_{\rm T}$, whereas for $\Upsilon$(2S), a slight increase in dissociation probability is observed with rising $p_{\rm T}$. Overall $p_{\rm T}$ dependence of $\Gamma_{\rm D}$ reflects the combined effects of the quarkonium decay width and its effective in-medium temperature. These contrasting behaviors offer important insights into the interplay between quarkonium binding energy and medium effects in heavy-ion collisions.
@@
Further, Fig. 3 and Fig. 4 demonstrate the impact of medium rotation on the $\Gamma_{\rm D}$, as a function of $p_{\rm T}$ for charmonium and bottomonium states, respectively. In the presence of a vortical background, the degeneracy in a quarkonium state is caused by spin-vorticity coupling. This implies that for $m_{j}=\pm 1$ state, the change in net $\Gamma_{\rm D}$ is prominent, which is an artifact of the term ($\sim m_{j}C/r^{2}$) appearing in the effective Hamiltonian. The spin–vorticity coupling alters the eigenenergies and radial wave functions in a manner that depends on the sign of $m_{j}$. The spin-vorticity coupling term ($\sim m_{j}C/r^{2}$) is expected to alter the short-distance ($r\rightarrow$ 0 ) behavior of quarkonium wave functions. At a large distance, the behavior of quarkonium wave functions is negligible. For the QCD vacuum potential, with $m_{j}=+1$, the additional term due to spin-orbit coupling lowers the effective potential barrier, thereby increasing the spatial extent of the wave function or the wave function becomes more peaked at short distances. Consequently, the dissociation probability rises, leading to a noticeably larger total decay width $\Gamma_{\rm D}$ compared to the non-rotating case. In contrast, for $m_{j}=-1$ carries the opposite sign in the rotational term, effectively making the potential slightly more confining. This suppresses the sensitivity of the quarkonium wave function at short distances. For the $m_{j}=0$ state, the quarkonium wave function is not affected due to the direct coupling to the rotation term, and naturally follows a trend similar to hydrogen atom wave functions. However, with the medium-modified color-singlet potential for quarkonia, the nature of the quarkonium wave function strongly depends on temperature and effective coupling constant. Further, the dissociation of the quarkonium state in-medium is led by the imaginary part of the potential, which induces the Landau damping. Therefore, the net decay width of a quarkonium state is a complex interplay between the temperature-dependent quarkonium wave function and its imaginary part of the complex potential. It is observed that the behavior of the quarkonium wave function at $m_{j}=-1$ and $m_{j}=0$ is almost similar. As a result, the net decay widths for the $m_{j}=0$ and $m_{j}=-1$ states remain closely aligned for both charmonium and bottomonium.
@@
Figure 5 illustrates $00^{\rm th}$-elements of the spin density matrix, $\rho_{00}$, as a function of charged particle multiplicity ($dN_{\rm ch}/d\eta>$) for $J/\psi$ (upper left), $\psi$(2S) (upper right), $\Upsilon(1S)$ (lower left), and $\Upsilon$(2S) (lower right) in Pb–Pb collisions at $\sqrt{s_{\rm NN}}$ = 5.02 TeV at mid-rapidity. We have obtained the charged particle multiplicity dependent $\rho_{00}$ for quarkonium state by averaging over $p_{\rm T}$ in the transverse momentum $p_{\rm T}$ range 1 $\leq p_{\rm T}\leq$ 30 GeV/c with the $p_{\rm T}$-distribution function $1/E_{\rm T}^{4}$ [31]. The $p_{\rm T}$-integrated charged particle multiplicity dependent spin alignment is given as,
@@
The results corresponding to observable $\rho_{00}$ are presented for various values of the circulation parameter $C$, which represents the strength of the rotation of the medium. In the absence of the experimental constraint on the $C$, we chose a set of $C$ values spreading from 0.01 fm to 5.0 fm. These chosen values represent the theoretical uncertainty for the prediction of $\rho_{00}$. The dashed horizontal line at $\rho_{00}=1/3$ corresponds to an unpolarized scenario and serves as the baseline for comparison. In the $C=0$ case, corresponding to an irrotational medium, all quarkonium states yield $\rho_{00}=1/3$ independent of charged particle multiplicity, as expected in the absence of spin-vorticity coupling. Finite values of $C$ induce a clear deviation from this baseline, indicating that medium rotation leads to spin-dependent dissociation and consequently generates spin alignment. Now at non-zero $C$, the $\rho_{00}$ for 1S states like $J/\psi$ and $\Upsilon$(1S), remains above $1/3$ over the entire multiplicity range for small and intermediate values of $C$, implying a relatively less dissociation of $m_{j}=0$ spin state. At low multiplicities, where the medium temperature is relatively smaller, the effect of vorticity is more pronounced and therefore, $\rho_{00}$ increases with $\langle dN_{\rm ch}/d\eta\rangle$. As the multiplicity increases, the rise in-medium temperature enhances the relative dissociation corresponding to $m_{j}=\pm 1$ degenerate states. While the effect of spin–vorticity coupling induced due to small values of C, remains marginal and becomes almost ineffective at higher temperatures or higher multiplicities. The combined dynamics of dissociation of $m_{j}=0,\text{and}\pm 1$ degenerate states, along with spin–vorticity coupling respective to small $C$ values, leads to a saturation $\rho_{00}$ at high multiplicities. For larger values of $C$, the enhanced splitting among the spin-dependent decay widths increases the dissociation of the $m_{j}=0$ state, causing $\rho_{00}$ to approach the unpolarized limit for $J/\psi$ and a change in spin alignment orientation for $\Upsilon$(1S) at high multiplicities.
@@
Figure 6 depicts the $p_{\rm T}$-dependent $\rho_{00}$ for $J/\psi$ (left panels) and $\psi$(2S) (right panels) for three multiplicities classes or centralities in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV. For the lowest multiplicity class (upper panels), $\rho_{00}$ exhibits a non-monotonic dependence on $p_{\rm T}$. At low transverse momentum, $\rho_{00}$ remains above the unpolarized limit, exhibiting a reduced dissociation of the $m_{j}=0$ state. With increasing $p_{\rm T}$, $\rho_{00}$ increases and reaches a maximum around $p_{\rm T}\simeq 15$ GeV/$c$, followed by a gradual decrease toward $\rho_{00}=1/3$ at higher $p_{\rm T}$. This behavior directly mirrors the $p_{\rm T}$ dependence of the net decay width $\Gamma_{\rm D}$, which is governed by the interplay between collisional damping, gluonic dissociation, and the momentum-dependent effective temperature. For higher multiplicity classes (middle and lower panels), the $p_{\rm T}$ dependence of $\rho_{00}$ develops a characteristic structure at low transverse momentum. In this region, $\rho_{00}$ initially decreases with increasing $p_{\rm T}$ up to $p_{\rm T}\sim 3$ GeV/$c$, followed by a rising trend at intermediate and high $p_{\rm T}$. This feature is an artifact of the $T_{\rm eff}$. Since $T_{\rm eff}$ enters explicitly into the imaginary part of the in-medium potential, this non-monotonic behavior is directly transmitted to $\Gamma_{\rm D}$ and consequently to $\rho_{00}$.
@@
In Fig. 7 spin alignment of $\Upsilon$(1S) and $\Upsilon$(2S) as a function of $p_{\rm T}$ are shown for chosen $C$ values and multiplicity classes in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV. At low multiplicity, $\rho_{00}$ for $\Upsilon$(1S) increases with $p_{\rm T}$, consistent with a reduction of the net decay width $\Gamma_{\rm D}$ as the effective temperature $T_{\rm eff}$ decreases at higher transverse momentum. At high multiplicities, $\rho_{00}$ decreases up to $p_{\rm T}\leq 10$ GeV, followed by a gradual increase toward $p_{\rm T}>10$ GeV. It reflects the non-monotonic $p_{\rm T}$ dependence of $T_{\rm eff}$, which enhances dissociation at low transverse momentum and reduces at higher $p_{\rm T}$. The systematic enhancement of $\rho_{00}$ with increasing $C$ for all multiplicity classes indicates that spin–vorticity coupling remains effective for the tightly bound $\Upsilon$(1S) state. In contrast, the $\Upsilon$(2S) state exhibits $\rho_{00}<1/3$ across the whole $p_{\rm T}$-range, and demonstarts the marginal dependence on both $p_{\rm T}$ and $C$. The smaller binding energy of the $\Upsilon$(2S) state enhances its sensitivity to thermal dissociation, which dominates over rotational effects and leads to a preferential dissociation of the $m_{j}=0$ component. Consequently, variations in $T_{\rm eff}$ primarily govern the observed $p_{\rm T}$ dependence for $\Upsilon$(2S), while spin-vorticity coupling introduces only marginal modifications.

### v2

@@ II.2.4 Gluonic Dissociation
The gluon-induced dissociation (or gluonic dissociation) of quarkonia in the QGP medium is another dissociation mechanism. This process arises when a thermal gluon interacts inelastically with a color-singlet quarkonium state, leading to a transition to a color-octet state, which subsequently breaks apart. The corresponding dissociation cross-section depends on the gluon energy spectrum and temperature profile of the medium. Following the formalism introduced by Peskin and Bhanot [49, 50] and extended to finite-temperature QCD [51, 52, 32, 33], the gluon-dissociation cross section for a quarkonium state is given by,
$\sigma_{{\rm d},nlm}(E_{g})=\frac{\pi^{2}\alpha_{s}^{u}E_{g}}{N_{c}^{2}}\sqrt{\frac{{m_{q}}}{E_{g}+E_{nlm}}}\\ \;\;\;\;\;\times\left(\frac{l|J_{nlm}^{q,l-1}|^{2}+(l+1)|J_{nlm}^{q,l+1}|^{2}}{2l+1}\right)$ (25)
where, $\alpha_{s}^{u}\simeq 0.76$ for bottom quark, and, $\alpha_{s}^{u}\simeq 0.87$ for charm quark. Here, $E_{g}$ is the incident gluon energy, and $E_{nlm}$ denotes the energy eigenvalue of the quarkonium state. The $g_{nlm}(r)$ and $h_{ql^{\prime}}(r)$ are the color-singlet and color-octet wave functions with orbital angular momenta $l$ and $l^{\prime}$, respectively. The octet wave function $h_{\rm ql^{\prime}}(r)$ has been obtained by solving the Schrödinger equation with the octet potential $V_{8}=\alpha_{\rm eff}/8r$. For a given value of $n$ and $l$, the color-octet wave function and the corresponding energy eigenvalue are identical for all three spin projection states $m$. The value of $q$ is determined by using the conservation of energy, $q=\sqrt{m_{q}(E_{g}+E_{nlm})}$.
The probability density $J_{nlm}^{ql^{\prime}}$ encodes the overlap between the initial bound and final continuum states and is computed using the in medium singlet and octet wave functions,
$J_{nlm}^{ql^{\prime}}=\int_{0}^{\infty}dr\;r\;g^{*}_{nlm}(r)\;h_{ql^{\prime}}(r)$ (26)
@@
This modified distribution plays an essential role in determining the gluonic dissociation rate and, consequently, the momentum-dependent dissociation of quarkonia in the QGP. The thermal averaged gluonic dissociation, $\Gamma_{{\rm gd},nlm}$, can then be obtained by folding the cross section with this modified Bose–Einstein distribution of thermal gluons in the particle rest frame [32, 33], as given by;
$\Gamma_{{\rm gd},nlm}(\tau,p_{\rm T},b)=\frac{g_{d}}{4\pi^{2}}\int_{0}^{\infty}\int_{0}^{\pi}\frac{dp_{g}\,d\theta\,\sin\theta\,p_{g}^{2}\sigma_{{\rm d},nlm}(E_{g})}{e^{\{\frac{\gamma E_{g}}{T_{\rm eff}}(1+v_{Q}\cos\theta)\}}-1}$ (28)
@@
Furthermore, we have obtained the net decay width $\Gamma_{\rm D}$ for quarkonia in-medium via taking the sum over Eq. (24), and (28), given as;
$\Gamma_{{\rm D},nlm}=\Gamma_{{\rm damp},nlm}+\Gamma_{{\rm gd},nlm}$ (29)
@@ II.3 Spin alignment of quarkonia
The dissociation probability of $m^{\rm th}$ spin state in the presence of rotation is given by;
${\mathcal{P}}_{m}=\exp\left[-\mathop{\text{\large$\int_{\text{\normalsize$\scriptstyle\kern-1.22911pt\tau_{0}$}}^{\text{\normalsize$\scriptstyle\tau_{f}$}}$}}\nolimits\Gamma_{{\rm D},nlm}(\tau,p_{\rm T})\;d\tau\right].$ (30)
Here $m$ takes on the values 0,+1,10,+1,-1, representing the three distinct spin projections corresponding to the states 0, ++1, and 1-1, respectively. The $\tau_{0}$ and $\tau_{f}$ are the proper times corresponding to the initial temperature $T_{0}$ and freeze-out temperature $T_{f}$ for the QGP medium. The dissociation rate, denoted as $\Gamma_{{\rm D},nlm}$ is intricately linked to both the rotational dynamics and the specific spin projection states, $m$.
The spin alignment of vector mesons is measured via the elements of the spin density matrix $\rho_{m,m^{\prime}}$, where $m$ and $m^{\prime}$ denote the spin projections along a chosen quantization axis. For a vector meson with spin 1, the spin density matrix is defined as;
$\rho_{m,m^{\prime}}=\frac{\langle A_{m}A^{*}_{m^{\prime}}\rangle}{\sum\limits^{k=+1}_{k=-1}\langle|A_{k}|^{2}\rangle}$ (31)
where $A_{m}$ denotes the production amplitude for the spin projection $m$, and the denominator ensures proper normalization. The density matrix satisfies the normalization condition
${\rm Tr(\rho)}=\sum\limits^{m=+1}_{m=-1}\rho_{m,m}=1$ (32)
The diagonal elements of the spin density matrix are given by
$\rho_{m,m}=\frac{\langle|A_{m}|^{2}\rangle}{\sum\limits^{k=+1}_{k=-1}\langle|A_{k}|^{2}\rangle}$ (33)
The production amplitude $A_{m}$ for the spin projection $m$, is related to the dissociation probability ${\mathcal{P}}_{m}$ via the relation
${\mathcal{P}}_{m}\equiv\langle|A_{m}|^{2}\rangle$ (34)
The $00^{\rm th}$ element of the $3\times 3$ spin density matrix is defined in terms of the $\Gamma_{{\rm D},nlm}$;
$\rho_{00}=\frac{{\mathcal{P}}_{0}}{{\mathcal{P}}_{+}+{\mathcal{P}}_{0}+{\mathcal{P}}_{-}}$ (35)
@@ III Results and Discussions
Quarkonium suppression and spin alignment provide complementary insights into the interaction between heavy quark–antiquark pairs and the quark–gluon plasma. Following that, we evaluated how different dissociation processes modify the spin alignment factor $\rho_{00}$ for the $J/\psi$, $\psi$(2S), $\Upsilon$(1S), and $\Upsilon$(2S) states, as defined in Eq. (35). Here, results are obtained by numerically solving the Schrödinger equation, which provides temperature, vorticity, and spin-projection-dependent eigenvalues and wave functions. The imaginary component of the potential, together with the quarkonium wave function, determines the collisional damping width, whereas the binding energy inferred from the eigenvalue governs the gluon-dissociation width. As reported in the earlier findings [35, 34], the collisional damping mechanism dominates over gluonic dissociation across the thermal evolution of the medium.
@@
The decay width illustrated in Fig. 2 shows the $p_{\rm T}$ dependence for $J/\psi$ (upper left), $\psi$(2S) (upper right), $\Upsilon$(1S) (lower left), and $\Upsilon$(2S) (lower right) in Pb–Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV in the absence in-medium rotation ($C=0$ fm) at $T$ = 175, 300, and 400 MeV. The results indicate a strong sensitivity of the dissociation rate to both temperature and momentum. At $T$ = 400 MeV, the $J/\psi$ exhibits an enhanced dissociation rate, which gradually decreases with increasing $p_{\rm T}$. The sudden drop in $\Gamma_{\rm D}$ around $p_{\rm T}\sim$ 10 GeV/c reflects that $T_{\rm eff}\ll T=400$ MeV. Further, at $T$ = 300, and 175 MeV, the effect of $T_{\rm eff}$ significantly diminishes and $\Gamma_{\rm D}$ decreases as a function of $p_{\rm T}$. A similar trend is evident for other quarkonium states, including $\psi$(2S), $\Upsilon$(1S), and $\Upsilon$(2S). In particular, at temperatures of 300 and 175 MeV, $\psi$(2S) exhibits a higher dissociation probability than $J/\psi$ across the entire $p_{\rm T}$ range. This can be attributed to the comparatively lower binding energy of $\psi$(2S) relative to $J/\psi$. Consequently, the reduced survival probability of $\psi$(2S) further supports the sequential melting scenario in the QGP medium. Furthermore, due to the higher mass of $\Upsilon$(1S) and $\Upsilon$(2S) states, which experience relatively stable effective temperatures as they traverse through the medium, as illustrated in Fig. 1. As a result, the dissociation probability for $\Upsilon$(1S) exhibits a modest decreasing trend towards higher $p_{\rm T}$. In contrast, for $\Upsilon$(2S), a slight increase in dissociation probability is observed with rising $p_{\rm T}$. Overall $p_{\rm T}$ dependence of $\Gamma_{\rm D}$ reflects the combined effects of the quarkonium decay width and its effective in-medium temperature. These contrasting behaviors offer important insights into the interplay between quarkonium binding energy and medium effects in heavy-ion collisions.
@@
Further, Fig. 3 and Fig. 4 demonstrate the impact of medium rotation on the $\Gamma_{\rm D}$, as a function of $p_{\rm T}$ for charmonium and bottomonium states, respectively. In the presence of a vortical background, the degeneracy in a quarkonium state is caused by spin-vorticity coupling. This implies that for $m_{j}=\pm 1$ state, the change in net $\Gamma_{\rm D}$ is prominent, which is an artifact of the term ($\sim m_{j}C/r^{2}$) appearing in the effective Hamiltonian. The spin–vorticity coupling alters the eigen energies and radial wave functions in a manner that depends on the sign of $m_{j}$. The spin-vorticity coupling term ($\sim m_{j}C/r^{2}$) is expected to alter the short-distance ($r\rightarrow$ 0 ) behavior of quarkonium wave functions. At a large distance, the behavior of quarkonium wave functions is negligible. For the QCD vacuum potential with $m_{j}=+1$, the additional term due to spin-orbit coupling lowers the effective potential barrier, thereby increasing the spatial extent of the wave function, or making it more peaked at short distances. Consequently, the dissociation probability rises, leading to a noticeably larger total decay width $\Gamma_{\rm D}$ compared to the non-rotating case. In contrast, for $m_{j}=-1$, the rotational term carries the opposite sign, effectively making the potential slightly more confining. This suppresses the sensitivity of the quarkonium wave function at short distances. For the $m_{j}=0$ state, the quarkonium wave function is not affected due to the direct coupling to the rotation term, and naturally follows a trend similar to hydrogen atom wave functions. However, with the medium-modified color-singlet potential for quarkonia, the nature of the quarkonium wave function strongly depends on temperature and effective coupling constant. Further, the dissociation of the quarkonium state in medium is led by the imaginary part of the potential, which induces the Landau damping. Therefore, the net decay width of a quarkonium state is a complex interplay between the temperature-dependent quarkonium wave function and its imaginary part of the complex potential. It is observed that the behavior of the quarkonium wave function at $m_{j}=-1$ and $m_{j}=0$ is almost similar. As a result, the net decay widths for the $m_{j}=0$ and $m_{j}=-1$ states remain closely aligned for both charmonium and bottomonium.
@@
Figure 5 illustrates $00^{\rm th}$-elements of the spin density matrix, $\rho_{00}$, as a function of charged particle multiplicity ($\langle dN_{\rm ch}/d\eta\rangle$) for $J/\psi$ (upper left), $\psi$(2S) (upper right), $\Upsilon$(1S) (lower left), and $\Upsilon$(2S) (lower right) in Pb–Pb collisions at $\sqrt{s_{\rm NN}}$ = 5.02 TeV at mid-rapidity. We have obtained the charged particle multiplicity dependent $\rho_{00}$ for quarkonium state by averaging over $p_{\rm T}$ in the transverse momentum $p_{\rm T}$ range 1 $\leq p_{\rm T}\leq$ 30 GeV/c with the $p_{\rm T}$-distribution function $1/E_{\rm T}^{4}$ [33]. The $p_{\rm T}$-integrated charged particle multiplicity dependent spin alignment is given as,
@@
The results corresponding to observable $\rho_{00}$ are presented for various values of the circulation parameter $C$, which represents the strength of the rotation of the medium. In the absence of the experimental constraint on the $C$, we chose a set of $C$ values spreading from 0.01 fm to 5.0 fm. These chosen values represent the theoretical uncertainty for the prediction of $\rho_{00}$. The dashed horizontal line at $\rho_{00}=1/3$ corresponds to an unpolarized scenario and serves as the baseline for comparison. In the $C=0$ case, corresponding to an irrotational medium, all quarkonium states yield $\rho_{00}=1/3$ independent of charged particle multiplicity, as expected in the absence of spin-vorticity coupling. Finite values of $C$ induce a clear deviation from this baseline, indicating the medium rotation leads to spin-dependent dissociation and consequently generates spin alignment. Now at non-zero $C$, the $\rho_{00}$ for 1S states like $J/\psi$ and $\Upsilon$(1S), remains above $1/3$ over the entire multiplicity range for small and intermediate values of $C$, implying a relatively less dissociation of $m_{j}=0$ spin state. At low multiplicities, where the medium temperature is relatively smaller, the effect of vorticity is more pronounced and therefore, $\rho_{00}$ increases with $\langle dN_{\rm ch}/d\eta\rangle$. As the multiplicity increases, the rise in-medium temperature enhances the relative dissociation corresponding to $m_{j}=\pm 1$ degenerate states. While the effect of spin–vorticity coupling induced due to small values of C, remains marginal and becomes almost ineffective at higher temperatures or higher multiplicities. The combined dynamics of dissociation of $m_{j}=0,\text{and}\pm 1$ degenerate states, along with spin–vorticity coupling respective to small $C$ values, leads to a saturation $\rho_{00}$ at high multiplicities. For larger values of $C$, the enhanced splitting among the spin-dependent decay widths increases the dissociation of the $m_{j}=0$ state, causing $\rho_{00}$ to approach the unpolarized limit for $J/\psi$ and a change in spin alignment orientation for $\Upsilon$(1S) at high multiplicities.
@@
Figure  6 depicts the $p_{\rm T}$-dependent $\rho_{00}$ for $J/\psi$ (left panels) and $\psi$(2S) (right panels) for three multiplicities classes or centralities in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV. For the lowest multiplicity class (upper panels), $\rho_{00}$ exhibits a non-monotonic dependence on $p_{\rm T}$. At low transverse momentum, $\rho_{00}$ remains above the unpolarized limit, exhibiting a reduced dissociation of the $m_{j}=0$ state. With increasing $p_{\rm T}$, $\rho_{00}$ increases and reaches a maximum around $p_{\rm T}\simeq 15$ GeV/$c$, followed by a gradual decrease toward $\rho_{00}=1/3$ at higher $p_{\rm T}$. This behavior directly mirrors the $p_{\rm T}$ dependence of the net decay width $\Gamma_{\rm D}$, which is governed by the interplay between collisional damping, gluonic dissociation, and the momentum-dependent effective temperature. For higher multiplicity classes (middle and lower panels), the $p_{\rm T}$ dependence of $\rho_{00}$ develops a characteristic structure at low transverse momentum. In this region, $\rho_{00}$ initially decreases with increasing $p_{\rm T}$ up to $p_{\rm T}\sim 3$ GeV/$c$, followed by a rising trend at intermediate and high $p_{\rm T}$. This feature is an artifact of the $T_{\rm eff}$. Since $T_{\rm eff}$ enters explicitly into the imaginary part of the in-medium potential, this non-monotonic behavior is directly transmitted to $\Gamma_{\rm D}$ and consequently to $\rho_{00}$.
@@
In Fig. 7 spin alignment of $\Upsilon$(1S) and $\Upsilon$(2S) as a function of $p_{\rm T}$ are shown for chosen $C$ values and multiplicity classes in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV. At low multiplicity, $\rho_{00}$ for $\Upsilon$(1S) increases with $p_{\rm T}$, consistent with a reduction of the net decay width $\Gamma_{\rm D}$ as the effective temperature $T_{\rm eff}$ decreases at higher transverse momentum. At high multiplicities, $\rho_{00}$ decreases up to $p_{\rm T}\leq 10$ GeV, followed by a gradual increase toward $p_{\rm T}>10$ GeV. It reflects the non-monotonic $p_{\rm T}$ dependence of $T_{\rm eff}$, which enhances dissociation at low transverse momentum and reduces at higher $p_{\rm T}$. The systematic enhancement of $\rho_{00}$ with increasing $C$ for all multiplicity classes indicates that spin–vorticity coupling remains effective for the tightly bound $\Upsilon$(1S) state. In contrast, the $\Upsilon$(2S) state exhibits $\rho_{00}<1/3$ across the whole $p_{\rm T}$-range, and demonstrates the marginal dependence on both $p_{\rm T}$ and $C$. The smaller binding energy of the $\Upsilon$(2S) state enhances its sensitivity to thermal dissociation, which dominates over rotational effects and leads to a preferential dissociation of the $m_{j}=0$ component. Consequently, variations in $T_{\rm eff}$ primarily govern the observed $p_{\rm T}$ dependence for $\Upsilon$(2S), while spin-vorticity coupling introduces only marginal modifications.